OpenStudy (anonymous):

Considering these two equations y = 2x + 2 and y = 1/2x + 5 What does the x equal in the solution of the system of equations

OpenStudy (anonymous):

just u have to solve the equations

OpenStudy (anonymous):

y = 2x + 2 and y = 1/2x + 5 2x + 2 = 1/2x + 5 SOLVE.

OpenStudy (anonymous):

check if x=2

OpenStudy (anonymous):

thanks people! x=2 is correct!

OpenStudy (anonymous):

LOL basically @matricked narrowed it down to like the answer

OpenStudy (anonymous):

yeah. but thanks for explaining it to me.

OpenStudy (anonymous):

y = 2x + 2 and y = 1/2x + 5 2x + 2 = 1/2x + 5 -2 -2 ---- ---- 2x = 1/2x + 3 -1/2x -1/2x ------ -------- 1.5x = 3 ---- ---- 1.5 1.5 x = 2

OpenStudy (anonymous):

OH. i get it now. thanks!

OpenStudy (anonymous):

yeah it was not that hard LOL :) yw :)